# Architecture independent generalization bounds for overparametrized deep ReLU networks

### v4

@@ 1.2.1. Remarks
For Theorem 1.3, it is not necessary for ${\mathcal{S}}^{test}_{W_{1}^{*}}$ and ${\mathcal{S}}^{train}_{W_{1}^{*}}$ to be identically distributed. We also note that in (1.24), we have that
@@
for all $\widetilde{x}\in{\mathcal{S}}^{test}$ and $x\in{\mathcal{S}}^{train}$, where $X^{(0)}$ is a matrix with columns formed by the $x^{(0)}$ vectors in $\mathcal{S}^{train}$, and $P_{X^{(0)}}=X^{(0)}(X^{(0)})^{+}$. In other words, only the component of $x$ in the linear span of the training data are relevant for the generalization error. Moreover, we point out that for all training data in the set
@@ 3.1. Interpretation of results
is the average training error for points in $\mathcal{S}^{train}$ which minimize the Chamfer distance to the test points. [^1] Here, $NN$ stands for nearest neighbor. Note that training points might contribute more than once to $\mathcal{E}_{\underline{\theta}_{*}}^{testNN}$. This proves the theorem. ∎

### v5

@@ 1.2. Generalization bound
In particular, if zero loss training is achieved, the trained weight $W_{1}^{*}$ acts projectively, and is supported on the span of the training data.
@@ 1.2.1. Remarks
For Theorem 1.3, it is not necessary for ${\mathcal{S}}^{test}_{W_{1}^{*}}$ and ${\mathcal{S}}^{train}_{W_{1}^{*}}$ to be identically distributed.
We also note that in (1.24), the trained weight $W_{1}^{*}$ acts projectively, and is supported on the span of the training data if zero training loss achieved. This means that
@@
for all $\widetilde{x}\in{\mathcal{S}}^{test}$ and $x\in{\mathcal{S}}^{train}$, where $X^{(0)}$ is a matrix with columns formed by the $x^{(0)}$ vectors in $\mathcal{S}^{train}$, and $P_{X^{(0)}}=X^{(0)}(X^{(0)})^{+}=P_{X^{(0)}}^{T}$ projects orthogonally onto ${\rm range}(X^{(0)})$. Here, $X_{0}^{+}$ is the generalized inverse, and $X_{0}^{+}=(X_{0}^{T}X_{0})^{-1}X_{0}^{T}$ when $X_{0}$ has full rank. In other words, only the component of $x$ in the linear span of the training data are relevant for the generalization error. Moreover, we point out that for all training data in the set
@@
This means that the trained network projects the test data set onto the span of the training data set, which equals ${\rm range}(X^{(0)})$ in ${\mathbb{R}}^{M_{0}}$. Therefore, the components of test data orthogonal to the span of the training data do not enter the generalization bound. We interpret this a key reason for the absence of overfitting.
@@ 3.1. Interpretation of results
is the average training error for points in $\mathcal{S}^{train}$ which minimize the Chamfer distance to the test points. [^1] Here, $NN$ stands for nearest neighbor. Note that training points might contribute more than once to $\mathcal{E}_{\underline{\theta}_{*}}^{testNN}$.
If zero training loss is achieved, then
$f_{\underline{\theta}_{*}}(x)=g_{\underline{\theta}_{*}}(W_{1}^{*}x+b_{1}^{*})=y\,$ (B.9)
for all $(x,y)\in{\mathcal{S}}^{train}$. This is equivalent to
$W_{1}^{*}x+b_{1}^{*}=g_{\underline{\theta}_{*}}^{-1}(y)$ (B.10)
where the right hand side denotes a (possibly non-unique) preimage of $y$ in the first layer. This implies that for the data matrix $X_{0}=[x_{1}\cdots x_{n}]$,
$W_{1}^{*}X_{0}=A:=[g_{\underline{\theta}_{*}}^{-1}(y_{1})-b_{1}^{*}\;\;\cdots\;\;g_{\underline{\theta}_{*}}^{-1}(y_{n})-b_{1}^{*}]$ (B.11)
and hence,
$W_{1}^{*}=AX_{0}^{+}\,.$ (B.12)
For the orthogonal projector $P_{X_{0}}=X_{0}X_{0}^{+}$ onto the range of $X_{0}$, we therefore find that $W_{1}^{*}P_{X_{0}}=AX_{0}^{+}(X_{0}X_{0}^{+})=AX_{0}^{+}=W_{1}^{*}$ acts projectively, and is supported on the range of $X_{0}$, which equals the span of the training data.
This proves the theorem. ∎